l Verizon 8:41 AM * 87%- chapter2 PDF-23 KB Close 3 of 3 Blue) is 610 or 60%, since there are 5 with blue stripes and 3 red for a total of 8 but 2 of red with blue stripes are counted twice and need to be subtracted from the total or looking at the chart you see there are 5 blue stripes and 1 additional red that you havent included. The complement would be Green n White which has 4 total. Now if we given information this is now called a conditional probability such as finding the probabil ity that the marble had a white stripe given it was green, P(x-WI G) is 4/7 since there are only 7 green and of them 4 had white stripes. For practice find Px-Red White),P(x-Green U Blue), P(x-WIR). Now assume you assume you are given a sample of 200 people and told that 60% of them voted For some particular issue. Of those that voted for 75% were men ( must have been sports related), and of those that voted No 60% were women. Make a contingency table and then find Px-Male No),P(x-Female U Yes), P(x MaleJNo) Email Save
Answer #1
Voted For Voted Against Total
Male 90 32 122
Female 30 48 78
Total 120 80 200

P(x=Male\cap No)=32/200=0.16

P(x=Female\cup Yes)=1-(32/200)=0.84

P(x=Male| No)=32/80=0.4